Periodic Inventory by Three Methods; Cost of Merchandise Sold The units of an item available for sale during the year were as follows: Jan. 1 Inventory 50 units @ $114 Mar. 10 Purchase 60 units @ $124 Aug. 30 Purchase 30 units @ $128 Dec. 12 Purchase 60 units @ $134 There are 80 units of the item in the physical inventory at December 31. The periodic inventory system is used. Determine the inventory cost and the cost of merchandise sold by three methods. Round interim calculations to one decimal and final answers to the nearest whole dollar.

Cost of Merchandise Inventory and Cost of Merchandise Sold
Inventory Method Merchandise Inventory Merchandise Sold
First-in, first-out (FIFO) $ $
Last-in, first-out (LIFO)
Weighted average cost
Answer #1
Unit Unit cost Total Cost
Beginning inventory 50 114 5700
Mar 10 Purchase 60 124 7440
Aug 30 Purchase 30 128 3840
Dec 12 Purchase 60 134 8040
Total 200 25020

Determine the inventory cost and the cost of merchandise sold by three methods. Round interim calculations to one decimal and final answers to the nearest whole dollar.

Cost of Merchandise Inventory and Cost of Merchandise Sold
Inventory Method Merchandise Inventory Merchandise Sold
First-in, first-out (FIFO) 60*134+20*128 = 10600 $25020-10600 = 14420
Last-in, first-out (LIFO) 50*114+30*124 = 9420 25020-9420 = 15600
Weighted average cost 25020/200*80 = 10008 25020-10008 = 15012